A Complete Bibliography of Publications in Communications in Statistics: Simulation

and Computation: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

.00001 [Mos78]. .01 [Mos78]. 2^{k−p} [NMC78]. 2 × 2 [Sub79]. _{u} [Win78]. C [Smi76]. χ^{2} [FS79]. F [Sub79]. k [CTW78]. L [Ode77a]. L_{1} [KGS77, SHC77, GHJ^{+}77]. s [Ode77b]. t [BOJ78, BO78, Mos78, MC76b, OJ77]. U [Wil78].

-distribution [MC76b, OJ77]. -sample [Smi76]. -statistic [Ode77b, Ode77a]. -statistics [Wil78]. -test [BOJ78, BO78]. -variates [Sub79].

absolute [AEH77, AF77, AFK79, BR77, BC77, Gen77, KG77, NW77, RC77]. adding [HSB79]. algorithm [AF77, AFK79, NW77]. Algorithms [BR77, Gun79, Mil78]. alternatives [FS79]. analyses [Mil78]. analysis [BOJ78, GHJ^{+}77, GS76, Kle77, LH77, Nag76]. Analytical [GA79]. any [Pre77]. application [LH77]. approach [GKI78, Gun79]. Approximate [Cro79, BS79, SRMS79, SHC77]. approximating [TW78]. approximation [Bad76, GHJ^{+}77, MC76b, RC77]. Approximations [IC78, Ury77]. area [OH77]. ARMA [GKI78, WG78]. armchair [Mil78]. aspects [Lor79]. asymmetric [Her79]. Asymptotic [AS78, DWB76]. attribute [Lor79].

barrier [Win76]. barrier-function [Win76]. based [Sub79, Wil78]. Bayes [Bro78]. Benard [BM76]. beta [Joh77]. between [Nag79]. bias [CP76]. biased [CH78, Gun79]. bivariate [GS78, Smi76]. block [Ode77c]. Board [Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d]. bound [Pre77].

calculation [Len76]. Carlo [Blo79, GWM79, Joh79, Kle77, LH78, Wil78]. case [BC77, Gup76]. cases [Smi76]. categories [HSB79]. category [HSB79]. cell [GS76, SD77]. censored [CP76]. central [TW78]. certainty [Bro78]. Charts [Mos78, Gue79]. chi [SRMS79]. chi-square [SRMS79]. choice [Bro78]. chosen [PJB79]. class [Che77]. classification [AEH77]. codes [GHJ^{+}77]. coefficient [GS78, MC76a]. combinations [PKK77]. combinatorial [Gol78]. combining [HSB79]. Comment [Tuk78]. Comments [Aka78, Gra78, HSB79, Rob78, GKM78]. common [MG79]. comparing [GHJ^{+}77]. Comparison [DKH77, GBE78, Mur76, Bro78, GS76, LN76, LH78, Smi76, Smi79, Ury77]. Comparisons [PJB79, GWM79]. complex [Gup76]. component [Nag76]. computation [BM76]. computational [Lor79]. computations [KG77, Pai78]. computer [KGS77, LH77]. computing [Woo77]. conditional [GBE78]. conditioned [OJ77]. Confidence [CTW78, Che77, Cro79, GBE78, MG79, OH77]. Consistency [And77]. control [Joh79, Ode77c]. Cornish [BS79, DH79]. Cornish-Fisher [DH79]. Correction [CP76]. correlated [CTW78]. correlation [GS78, MC76a, Nag76, OJ77]. cost [NMC78]. cost-optimal [NMC78]. covariance [DS77, Joh77, Ric76]. critical [IC78, Joh77, Mon78]. cumulative [Len76]. curve [AFK79, DH79, Wei78]. curves [PJB79].

data [SD77, Sun76, WG78]. decomposition [FW78]. density [Blo79]. dependent [Cro79, CM79, LN76]. design [AS78, Kle77]. designed [TM78]. designs [Ode77c]. determinant [Nag76]. Determining [Win78]. deviation [AFK79]. difference [Joh77]. different [Nag79, PJB79]. direct [BTT78]. discrete [DWB76, GHJ^{+}77]. Distribution [Her79, Bad76, CP76, DH79, DB77, GBE78, Gup76, LCO79, MC76a, MC76b, Nag76, Ode77a, Ode77b, OJ77, OH77, RC77, Ury77, WP77]. distributions [DWB76, FW78, Ode77c, PKK77, PJB79, SRMS79, TW78, Woo77]. diversity [HSB79]. double [GBE78]. doubly [MC76b]. duality [And77]. dummy [AF77].

Editorial [Kru76, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d]. efficiency [DWB76]. eight [PJB79]. Elteren [BM76]. empirical [LN76, Smi76, Smi79]. equality [LN76]. equations [And77, Sun76]. equiprobability [SRMS79]. equivalence [Bro78]. error [BOJ78, DH79, KG77]. errors [NW77]. estimate [Blo79]. estimates [AEH77, BTT78, RC77]. estimation [AS78, BR77, Gen77, Gol78, SD77, Wil78, Win76]. Estimator [SHC77, DB77]. estimators [AF77, CH78, CM79, CP76, DKH77, Woo77]. evaluating [VW76]. Evaluation [MG79, LH78, WP77]. Exact [SRMS79, CI78, Len76, Ode77a]. Examining [KG77]. example [Joh79]. expansion [DH79]. experiment [Kle77]. experiments [Joh79]. exponential [GBE78, PKK77, Sun76]. Extended [Ode77b, Ode77c].

F [Len76]. factorials [NMC78]. families [Sun76]. finding [Mon78, Mos78]. first [PJB79]. Fisher [BS79, DH79, FW78, MC76a]. fit [FS79, LCO79]. fitting [AFK79, DH79, Wei78]. five [CH78]. fixed [Che77]. fixed-width [Che77]. four [PJB79]. Fourier [Blo79, WP77]. Fox [Gue79]. fractional [NMC78]. frequency [PJB79]. Friedman [Ode77b]. function [Bad76, Pre77, Win76]. functions [Ric76, WP77, Woo77]. Further [BS79].

gamma [DB77, TW78, Woo77]. general [Pai78]. generalized [Wil78]. generated [Nag79]. generating [KGS77]. geodetic [Siv78]. global [Gol78]. Goodness [LCO79, FS79]. Goodness-of-fit [LCO79]. graph [LH77]. grouped [SD77].

hazard [GA79]. homogeneity [GWM79, KGC79].

II [MQ79]. illustrating [Kle77]. incomplete [Sun76]. independent [Sub79]. indices [HSB79]. integral [Blo79]. interactive [VW76]. intervals [Che77, CTW78, Cro79, GBE78, MG79]. introduction [Gen77]. iterative [Sun76].

Johnson [Win78].

large [Gol78]. largest [CTW78]. layout [Ode77a, Ode77b]. Least [BC77, Gen77, AEH77, AF77, BR77, DW78, KG77, RC77]. length [FW78]. likelihood [CP76, Siv78, Sun76]. limits [OH77]. line [CH78]. linear [AS78, And77, AF77, BC77, GHJ^{+}77, Pai78, PKK77, Pre77]. lines [Smi79]. locally [Mon78]. location [Siv78]. lognormal [Win76].

Mann [Ury77]. MANOVA [Gup76]. Markov [Cro79, CM79]. matrices [Joh77]. matrix [DS77, Nag76, Ric76]. Maximum [Siv78, CP76]. mean [Che77, CTW78, Joh77, MG79]. means [Her79, Mur76]. measures [NG79]. method [Joh79, KGS77, MG79, Sun76]. Methodology [GHJ^{+}77]. methods [LN76, LH78, Nag79, Wei78, Win76]. midpoint [SD77]. Minicommunications [Goo76, GSJ76, GSK76]. minimization [BC77]. minimum [NW77, Woo77]. Mises [FW78]. model [AEH77, AF77, DH79]. modeling [GKI78]. models [Pai78, Win76, WG78]. moment [DB77]. moments [PJB79, Ric76]. Monte [Blo79, GWM79, Joh79, Kle77, LH78, Wil78]. Monte-Carlo [GWM79]. mortality [GA79]. Moving [Win76]. multiple [Kle77, LH78, Lor79]. multivariate [Joh77, Mon78, TW78].

noise [FS79, Joh79]. noncentral [MC76b, TW78]. nonlinear [DW78]. Nonnull [Gup76]. nonparametric [BM76, LH78, TM78]. normal [And77, Bad76, Che77, CTW78, CP76, DS77, DKH77, GS78, Geh79, MG79, OJ77, OH77, Woo77]. normed [Pre77]. note [Gue79, Mon78]. Numerically [Pai78].

observation [Siv78]. Obtaining [AEH77, AF77]. odds [Sub79]. Odeh [Gue79]. one [BOJ78, BO78, OJ77]. Optimal [BTT78, Bro78, Mon78, NMC78, Wei78]. Optimization [Rus78, Vis78]. optimum [Gol78, Siv78]. order [DS77, Nag79, PKK77]. oriented [KGS77]. orthant [Geh79]. overview [Rus78].

packages [VW76]. Page [Ode77a]. parallelism [Smi79]. Parameter [NG79, DB77, LCO79, Win76]. parameters [Win78, Woo77]. partitioning [BO78]. Pearson [BS79, DH79]. percentage [BS79, Mos78, PKK77, PJB79]. percentiles [DKH77]. performance [GH78]. Performing [Mil78]. piecewise [BC77]. pilot [Wil78]. plus [FS79]. Point [Gol78, DKH77]. points [BS79, Mos78, PKK77, PJB79]. population [GS78]. populations [CTW78, MG79]. portfolio [Bro78]. portfolios [Bro78]. positive [Geh79]. Posterior [Ric76]. Power [MQ79, BO78, BTT78, FS79]. present [IC78]. principal [Nag76]. probabilities [Geh79, Len76]. probability [WP77]. problem [AFK79, Bro78, LH77, Smi76, Vis78]. problems [Gol78, Gue79, KGS77]. procedure [Kle77, TM78]. procedures [GH78, LH78, Lor79, Wil78]. process [Vis78]. program [VW76]. programming [AF77, Gun79]. programs [GS76]. proportion [Cro79, CM79]. proportions [LN76]. purpose [AF77].

quadrivariate [Geh79]. Quantile [SD77].

randomized [Ode77c]. range [Mos78]. rank [Ode77c, Smi76]. ranked [Che77]. ranking [Kle77]. ranks [CI78, DWB76]. ratio [Sub79]. ratios [PKK77]. reduces [Joh79]. region [IC78]. regression [BC77, CH78, DW78, Gun79, KGS77, KG77, NW77, RC77, Smi79]. Rejoinder [GKM78]. relations [Nag79]. reliability [GA79]. representation [Geh79]. representations [GA79]. residual [LH77]. residuals [Pre77, RC77]. restricted [BR77]. resultant [FW78]. revised [AFK79]. rho [IC78]. right [CP76]. right-censored [CP76]. robust [DW78]. Robustness [Kle77]. rounding [KG77]. roundoff [Mil78].

S [Win78]. same [PJB79]. sample [BOJ78, BO78, GS78, GWM79, Gue79, Mur76, Smi76]. samples [FW78, PKK77]. sampling [Lor79, RC77]. scalar [Ric76]. Selection [NMC78, GH78]. sequential [Wil78]. series [WP77]. setting [MG79]. several [GS76, Joh77, LN76, Smi79]. shape [DB77]. signal [FS79]. significance [Sub79]. simple [MC76b, RC77]. simplex [AFK79]. simulation [CH78, GA79, Mur76, TM78]. sites [Siv78]. size [GWM79, GS76, Gue79]. skewness [NG79]. Small [GWM79]. smoothing [BTT78]. solution [Gue79, Sun76]. Some [CI78, Nag79, GH78, LH78, Mur76, Ury77]. spatial [Siv78]. Spearman [IC78]. special [AF77, BC77]. spectrum [BTT78]. splines [AS78]. square [SRMS79]. squared [CI78, DWB76]. squares [DW78]. stable [Pai78]. statistic [BM76, Gup76, Mon78, Ode77a, Ode77b, SRMS79]. Statistical [Wei78, Mil78, VW76]. statistics [DS77, Nag79, Ode77c, PKK77, Rus78, Wil78]. storage [LH77]. straight [CH78]. Student [BOJ78, BO78, Mos78]. studies [GA79, MQ79, TM78]. study [Blo79, CH78, Wil78]. Subroutines [Woo77]. subset [GH78]. sum [DWB76, NW77]. sunspot [WG78]. surveying [Siv78]. systems [PJB79].

Table [PKK77, Sub79]. Tables [OH77, CI78, Ode77b, Ode77c]. tail [OH77]. Techniques [DW78, Kle77]. test [BOJ78, BO78, CI78, DWB76, FS79, KGS77, Mon78]. testing [Joh77, LN76]. Tests [KGC79, Sub79, GWM79, LCO79, MQ79, Mur76, Smi76, Smi79]. theory [LH77]. three [Win76]. three-parameter [Win76]. ties [IC78]. trace [Joh77]. transformation [MC76a]. treatment [Ode77c]. trials [Cro79, CM79]. trimmed [Her79]. truncated [GS78]. truncations [Win76]. two [AEH77, HSB79, Joh77, LCO79, MG79, Mur76, Ode77a, Ode77b, Smi76, Sub79, Wil78]. two-way [AEH77, Ode77a, Ode77b]. type [DH79, BOJ78]. type-VI [DH79]. types [Smi76].

Unbiased [Joh77, Woo77]. unconditional [GBE78]. unequal [GS76]. uniformity [MQ79]. univariate [Pai78, Smi76]. upper [Joh77, Mos78, Pre77]. use [DH79, GA79, WP77]. useful [FW78, Nag76]. Using [SHC77, AS78, Gue79].

Validation [CM79]. value [AEH77, AF77, BR77, BC77]. values [Gen77, Joh77, KG77, Mon78]. variables [AF77]. variance [GS76, KGC79, Woo77]. variances [GWM79]. variate [Joh79]. variates [Sub79]. vectors [Joh77]. versus [Ode77c]. VI [DH79]. via [And77]. way [AEH77, Ode77a, Ode77b]. Weibull [LCO79]. weighted [NW77]. Whitney [Ury77]. width [Che77]. Wiener [Vis78]. Wilcoxon [Ury77]. Wilks [Gup76]. without [IC78]. Wolfer [WG78].

References

Armstrong:1977:OLA

[AEH77] Ronald D. Armstrong, Joyce J. Elam, and John W. Hultz. Obtaining least absolute value estimates for a two-way classification model. Communications in Statistics: Simulation and Computation, 6(4): 365–381, 1977. CODEN CSSCDB. ISSN 0361-0918.

Armstrong:1977:SPL

[AF77] Ronald D. Armstrong and Edward Frome. A special purpose linear programming algorithm for obtaining least absolute value estimators in a linear model with dummy variables. Communications in Statistics: Simulation and Computation, 6(4):383–398, 1977. CO-DEN CSSCDB. ISSN 0361-0918.

Armstrong:1979:RSA

[AFK79] Ronald D. Armstrong, E. L. Frome, and D. S. Kung. A revised simplex algorithm for the absolute deviation curve fitting problem. Communications in Statistics: Simulation and Computation, 8(2): 175–190, 1979. CODEN CSSCDB. ISSN 0361-0918.

Akaike:1978:C

[Aka78] Hirotugu Akaike. Comments. Communications in Statistics: Simulation and Computation, 7(1):84–87, 1978. CODEN CSSCDB. ISSN 0361-0918.

Andriano:1977:CNE

[And77] Kim Andriano. Consistency of the normal equations via linear duality. Communications in Statistics: Simulation and Computation, 6 (1):97–99, 1977. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1977:EBa

[Ano77a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 6(1):ebi–ebii, 1977. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1977:EBb

[Ano77b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 6(2):ebi–ebii, 1977. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1977:EBc

[Ano77c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 6(3):ebi–ebii, 1977. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1977:EBd

[Ano77d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 6(4):ebi–ebii, 1977. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1978:EBa

[Ano78a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 7(1):ebi–ebii, 1978. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1978:EBb

[Ano78b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 7(3):ebi–ebii, 1978. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1978:EBc

[Ano78c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 7(4):ebi–ebii, 1978. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1978:EBd

[Ano78d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 7(5):ebi–ebii, 1978. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1979:EBa

[Ano79a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 8(1):ebi–ebii, 1979. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1979:EBb

[Ano79b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 8(2):ebi–ebii, 1979. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1979:EBc

[Ano79c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 8(3):ebi–ebii, 1979. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1979:EBd

[Ano79d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 8(4):ebi–ebii, 1979. CODEN CSSCDB. ISSN 0361-0918.

Agarwal:1978:ADE

[AS78] Girdhar G. Agarwal and W. J. Studden. Asymptotic design and estimation using linear splines. Communications in Statistics: Simulation and Computation, 7(4):309–319, 1978. CODEN CSSCDB. ISSN 0361-0918.

Badhe:1976:NAN

[Bad76] Sahadeo K. Badhe. New approximation of the normal distribution function. Communications in Statistics: Simulation and Computation, 5(4):173–176, 1976. CODEN CSSCDB. ISSN 0361-0918.

Bartels:1977:LAV

[BC77] Richard H. Bartels and Andrew R. Conn. Least absolute value regression: a special case of piecewise linear minimization. Communications in Statistics: Simulation and Computation, 6(4):329–339, 1977. CODEN CSSCDB. ISSN 0361-0918.

Bloxom:1979:FID

[Blo79] Bruce Bloxom. A Fourier integral density estimate: a Monte Carlo study. Communications in Statistics: Simulation and Computation, 8(4):391–396, 1979. CODEN CSSCDB. ISSN 0361-0918.

Brunden:1976:BVE

[BM76] M. N. Brunden and N. R. Mohberg. The Benard–Van Elteren statistic and nonparametric computation. Communications in Statistics: Simulation and Computation, 5(4):155–162, 1976. CODEN CSS-CDB. ISSN 0361-0918.

Browne:1978:PPO

[BO78] Richard H. Browne and D. B. Owen. A partitioning of the power in the one sample Student t-test. Communications in Statistics: Simulation and Computation, 7(6):605–617, 1978. CODEN CSSCDB. ISSN 0361-0918.

Browne:1978:ATE

[BOJ78] Richard H. Browne, D. B. Owen, and Faming Ju. An analysis of the Type I error in the one sample Student t-test. Communications in Statistics: Simulation and Computation, 7(6):593–603, 1978. CO-DEN CSSCDB. ISSN 0361-0918.

Barrodale:1977:ARL

[BR77] I. Barrodale and F. D. K. Roberts. Algorithms for restricted least absolute value estimation. Communications in Statistics: Simulation and Computation, 6(4):353–363, 1977. CODEN CSSCDB. ISSN 0361-0918.

Brown:1978:PCP

[Bro78] Stephen J. Brown. The portfolio choice problem: comparison of certainty equivalence and optimal Bayes portfolios. Communications in Statistics: Simulation and Computation, 7(4):321–334, 1978. CO-DEN CSSCDB. ISSN 0361-0918.

Bowman:1979:FAP

[BS79] K. O. Bowman and L. R. Shenton. Further approximate Pearson percentage points and Cornish–Fisher. Communications in Statistics: Simulation and Computation, 8(3):231–244, 1979. CODEN CSSCDB. ISSN 0361-0918.

Byrd:1978:OSD

[BTT78] R. H. Byrd, R. A. Tapia, and J. R. Thompson. Optimal smoothing of direct estimates of the power spectrum. Communications in Statistics: Simulation and Computation, 7(4):335–344, 1978. CO-DEN CSSCDB. ISSN 0361-0918.

Carmer:1978:SSF

[CH78] S. G. Carmer and W. T. Hsieh. A simulation study of five biased estimators for straight line regression. Communications in Statistics: Simulation and Computation, 7(6):529–548, 1978. CODEN CSSCDB. ISSN 0361-0918.

Chen:1977:CFW

[Che77] Hubert J. Chen. A class of fixed-width confidence intervals for a ranked normal mean. Communications in Statistics: Simulation and Computation, 6(2):137–156, 1977. CODEN CSSCDB. ISSN 0361-

Conover:1978:SET

[CI78] W. J. Conover and Ronald L. Iman. Some exact tables for the squared ranks test. Communications in Statistics: Simulation and Computation, 7(5):491–513, 1978. CODEN CSSCDB. ISSN 0361-0918.

Crow:1979:VEP

[CM79] Edwin L. Crow and Martin J. Miles. Validation of estimators of a proportion from Markov dependent trials. Communications in Statistics: Simulation and Computation, 8(1):25–52, 1979. CODEN CSSCDB. ISSN 0361-0918.

Custer:1976:CBM

[CP76] S. W. Custer and S. H. Pam. Correction for bias in maximum likelihood estimators of in a right-censored normal distribution. Communications in Statistics: Simulation and Computation, 5(1):15–22, 1976. CODEN CSSCDB. ISSN 0361-0918.

Crow:1979:ACI

[Cro79] Edwin L. Crow. Approximate confidence intervals for a proportion from Markov dependent trials. Communications in Statistics: Simulation and Computation, 8(1):1–24, 1979. CODEN CSSCDB. ISSN 0361-0918.

Chen:1978:CIL

[CTW78] Hubert J. Chen, Paul J. Tsai, and Mike W. Wang. Confidence intervals for the largest mean from k correlated normal populations. Communications in Statistics: Simulation and Computation, 7(3): 205–222, 1978. CODEN CSSCDB. ISSN 0361-0918.

Dusenberry:1977:MES

[DB77] W. E. Dusenberry and K. O. Bowman. The moment estimator for the shape parameter of the gamma distribution. Communications in Statistics: Simulation and Computation, 6(1):1–19, 1977. CODEN CSSCDB. ISSN 0361-0918.

Davenport:1979:UCF

[DH79] James M. Davenport and Thomas A. Herring. On the use of curve fitting to model the error of the Cornish-Fisher expansion of the Pearson type-VI distribution the error of the Cornish-Fisher expansion of the Pearson type-VI distribution. Communications in Statistics: Simulation and Computation, 8(4):311–333, 1979. CODEN CSS-CDB. ISSN 0361-0918.

Dyer:1977:CPE

[DKH77] D. D. Dyer, J. P. Keating, and O. L. Hensley. Comparison of point estimators of normal percentiles. Communications in Statistics: Simulation and Computation, 6(3):269–283, 1977. CODEN CSSCDB. ISSN 0361-0918.

Davis:1977:CMN

[DS77] C. S. Davis and M. A. Stephens. The covariance matrix of normal order statistics. Communications in Statistics: Simulation and Computation, 6(1):75–81, 1977. CODEN CSSCDB. ISSN 0361-0918.

Dennis:1978:TNL

[DW78] John E. Dennis, Jr. and Roy E. Welsch. Techniques for nonlinear least squares and robust regression. Communications in Statistics: Simulation and Computation, 7(4):345–359, 1978. CODEN CSSCDB. ISSN 0361-0918.

Duran:1976:AES

[DWB76] B. S. Duran, M. Whiteside, and T. L. Boullion. Asymptotic efficiency of the sum of squared ranks test for discrete distributions. Communications in Statistics: Simulation and Computation, 5(1):33–42, 1976. CODEN CSSCDB. ISSN 0361-0918.

Francisco:1979:PGF

[FS79] Lee D. Kaiser Francisco and J. Samaniego. On the power of the χ^{2} goodness of fit test at signal plus noise alternatives. Communications in Statistics: Simulation and Computation, 8(1):75–90, 1979. CODEN CSSCDB. ISSN 0361-0918.

Fisher:1978:UDR

[FW78] N. I. Fisher and M. E. Willcox. A useful decomposition of the resultant length for samples from von Mises–Fisher distributions. Communications in Statistics: Simulation and Computation, 7(3):257–267, 1978. CODEN CSSCDB. ISSN 0361-0918.

Gaver:1979:AHR

[GA79] D. P. Gaver and M. Acar. Analytical hazard representations for use in reliability, mortality, and simulation studies. Communications in Statistics: Simulation and Computation, 8(2):91–111, 1979. CODEN CSSCDB. ISSN 0361-0918.

Grice:1978:CCU

[GBE78] John V. Grice, Lee J. Bain, and Max Engelhardt. Comparison of conditional and unconditional confidence intervals for the double exponential distribution. Communications in Statistics: Simulation and Computation, 7(5):515–524, 1978. CODEN CSSCDB. ISSN 0361-0918.

Gehrlein:1979:RQN

[Geh79] William V. Gehrlein. A representation for quadrivariate normal positive orthant probabilities. Communications in Statistics: Simulation and Computation, 8(4):349–358, 1979. CODEN CSSCDB. ISSN 0361-0918.

Gentle:1977:LAV

[Gen77] James E. Gentle. Least absolute values estimation: an introduction. Communications in Statistics: Simulation and Computation, 6(4): 313–328, 1977. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1978:PSS

[GH78] Shanti S. Gupta and Jason C. Hsu. On the performance of some subset selection procedures. Communications in Statistics: Simulation and Computation, 7(6):561–591, 1978. CODEN CSSCDB. ISSN 0361-0918.

Gilsinn:1977:MAC

[GHJ^{+}77] J. Gilsinn, K. Hoffman, R. H. F. Jackson, E. Leyendecker, P. Saunders, and D. Shier. Methodology and analysis for comparing discrete linear l_{1} approximation codes. Communications in Statistics: Simulation and Computation, 6(4):399–413, 1977. CODEN CSSCDB. ISSN 0361-0918.

Gray:1978:NAA

[GKI78] H. L. Gray, G. D. Kelley, and D. D. Mc Intire. A new approach to ARMA modeling. Communications in Statistics: Simulation and Computation, 7(1):1–77, 1978. CODEN CSSCDB. ISSN 0361-0918.

Gray:1978:RC

[GKM78] H. L. Gray, G. D. Kelley, and D. D. McIntire. Rejoinder to comments. Communications in Statistics: Simulation and Computation, 7(1): 92–95, 1978. CODEN CSSCDB. ISSN 0361-0918.

Golden:1978:PEG

[Gol78] Bruce L. Golden. Point estimation of a global optimum for large combinatorial problems. Communications in Statistics: Simulation and Computation, 7(4):361–367, 1978. CODEN CSSCDB. ISSN 0361-0918.

Good:1976:Ma

[Goo76] I. J. Good. Minicommunications. Communications in Statistics: Simulation and Computation, 5(1):81–84, 1976. CODEN CSSCDB. ISSN 0361-0918.

Granger:1978:C

[Gra78] C. W. J. Granger. Comments. Communications in Statistics: Simulation and Computation, 7(1):89–92, 1978. CODEN CSSCDB. ISSN 0361-0918.

Golhar:1976:CSA

[GS76] M. B. Golhar and J. H. Skillings. A comparison of several analysis of variance programs with unequal cell size. Communications in Statistics: Simulation and Computation, 5(1):43–53, 1976. CODEN CSSCDB. ISSN 0361-0918.

Gajjar:1978:SCC

[GS78] A. V. Gajjar and Kocherlakota Subrahmaniam. On the sample correlation coefficient in the truncated bivariate normal population. Communications in Statistics: Simulation and Computation, 7(5):455– 477, 1978. CODEN CSSCDB. ISSN 0361-0918.

Good:1976:Mb

[GSJ76] I. J. Good, Eugene Seneta, and Donald R. Jensen. Minicommunications. Communications in Statistics: Simulation and Computation, 5(2–3):141–147, 1976. CODEN CSSCDB. ISSN 0361-0918.

Good:1976:Mc

[GSK76] I. J. Good, S. R. Searle, and A. I. Khuri. Minicommunications. Communications in Statistics: Simulation and Computation, 5(4): 207–211, 1976. CODEN CSSCDB. ISSN 0361-0918.

Guenther:1979:NSS

[Gue79] William C. Guenther. A note on the solution of sample size problems using the odeh and Fox charts. Communications in Statistics: Simulation and Computation, 8(2):169–171, 1979. CODEN CSSCDB. ISSN 0361-0918.

Gunst:1979:APB

[Gun79] Richard F. Gunst. An approach to the programming of biased regression algorithms. Communications in Statistics: Simulation and Computation, 8(2):151–159, 1979. CODEN CSSCDB. ISSN 0361-

Gupta:1976:NDW

[Gup76] A. K. Gupta. Nonnull distribution of Wilks’ statistic for MANOVA in the complex case. Communications in Statistics: Simulation and Computation, 5(4):177–188, 1976. CODEN CSSCDB. ISSN 0361-0918.

Geng:1979:SSS

[GWM79] S. Geng, W. J. Wang, and C. Miller. Small sample size comparisons of tests for homogeneity of variances by Monte-Carlo. Communications in Statistics: Simulation and Computation, 8(4):379–389, 1979. CODEN CSSCDB. ISSN 0361-0918.

Hertsgaard:1979:DAT

[Her79] Doris Hertsgaard. Distribution of asymmetric trimmed means. Communications in Statistics: Simulation and Computation, 8(4):359– 367, 1979. CODEN CSSCDB. ISSN 0361-0918.

Hutcheson:1979:CDI

[HSB79] K. Hutcheson, L. R. Shenton, and R. C. Bailey. Comments on diversity indices when adding a category and combining two categories. Communications in Statistics: Simulation and Computation, 8(4): 399–401, 1979. CODEN CSSCDB. ISSN 0361-0918.

Iman:1978:ACR

[IC78] Ronald L. Iman and W. J. Conover. Approximations of the critical region for Spearman’s rho with and without ties present. Communications in Statistics: Simulation and Computation, 7(3):269–282, 1978. CODEN CSSCDB. ISSN 0361-0918.

John:1977:UUC

[Joh77] S. John. Unbiased and upper critical values of mean trace of multivariate beta for testing difference of two covariance matrices or several mean vectors. Communications in Statistics: Simulation and Computation, 6(1):89–96, 1977. CODEN CSSCDB. ISSN 0361-0918.

Johannes:1979:EHC

[Joh79] James M. Johannes. An example of how the control variate method reduces noise in Monte Carlo experiments. Communications in Statistics: Simulation and Computation, 8(4):335–347, 1979. CO-DEN CSSCDB. ISSN 0361-0918.

Kennedy:1977:ERE

[KG77] W. J. Kennedy and James E. Gentle. Examining rounding error in least absolute values regression computations. Communications in Statistics: Simulation and Computation, 6(4):415–420, 1977. CO-DEN CSSCDB. ISSN 0361-0918.

Keselman:1979:THV

[KGC79] H. J. Keselman, Paul A. Gaines, and Jennifer J. Clinch. Tests for homogeneity of variance. Communications in Statistics: Simulation and Computation, 8(2):113–129, 1979. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1977:COM

[KGS77] W. J. Kennedy, J. E. Gentle, and V. A. Sposito. A computer oriented method for generating test problems for L_{1} regression. Communications in Statistics: Simulation and Computation, 6(1):21–27, 1977. CODEN CSSCDB. ISSN 0361-0918.

Kleijnen:1977:RMR

[Kle77] Jack P. C. Kleijnen. Robustness of a multiple ranking procedure: a Monte Carlo experiment illustrating design and analysis techniques. Communications in Statistics: Simulation and Computation, 6(3): 235–262, 1977. CODEN CSSCDB. ISSN 0361-0918.

Krutchkoff:1976:E

[Kru76] Richard G. Krutchkoff. Editorial. Communications in Statistics: Simulation and Computation, 5(1):i–iii, 1976. CODEN CSSCDB. ISSN 0361-0918.

Littell:1979:GFT

[LCO79] Ramon C. Littell, James T. Mc Clave, and Walter W. Offen. Goodness-of-fit tests for the two parameter Weibull distribution. Communications in Statistics: Simulation and Computation, 8(3): 257–269, 1979. CODEN CSSCDB. ISSN 0361-0918.

Lentner:1976:ECC

[Len76] Marvin Lentner. On the exact calculation of cumulative f probabilities. Communications in Statistics: Simulation and Computation, 5 (4):149–154, 1976. CODEN CSSCDB. ISSN 0361-0918.

Lyons:1977:AGT

[LH77] N. I. Lyons and K. Hutcheson. An application of graph theory to a computer storage problem in residual analysis. Communications in Statistics: Simulation and Computation, 6(2):157–166, 1977. CO-DEN CSSCDB. ISSN 0361-0918.

Lin:1978:ESN

[LH78] F. A. Lin and J. K. Haseman. An evaluation of some nonparametric multiple comparison procedures by Monte Carlo methods. Communications in Statistics: Simulation and Computation, 7(2):117–128, 1978. CODEN CSSCDB. ISSN 0361-0918.

Levy:1976:ECS

[LN76] Kenneth J. Levy and Subhash C. Narula. An empirical comparison of several methods for testing the equality of dependent proportions. Communications in Statistics: Simulation and Computation, 5(4): 189–195, 1976. CODEN CSSCDB. ISSN 0361-0918.

Lorenzen:1979:CAM

[Lor79] Thomas J. Lorenzen. The computational aspects of multiple attribute sampling procedures. Communications in Statistics: Simulation and Computation, 8(4):291–309, 1979. CODEN CSSCDB. ISSN 0361-0918.

Mudholkar:1976:DFT

[MC76a] Govind S. Mudholkar and Yogendra P. Chaubey. On the distribution of Fisher’s transformation of the correlation coefficient. Communications in Statistics: Simulation and Computation, 5(4):163–172, 1976. CODEN CSSCDB. ISSN 0361-0918.

Mudholkar:1976:SAD

[MC76b] Govind S. Mudholkar and Yogendra P. Chaubey. A simple approximation for the doubly noncentral t-distribution. Communications in Statistics: Simulation and Computation, 5(2–3):85–92, 1976. CO-DEN CSSCDB. ISSN 0361-0918.

Maric:1979:EMS

[MG79] Nebojsa Marić and Franklin A. Graybill. Evaluation of a method for setting confidence intervals on the common mean of two normal populations. Communications in Statistics: Simulation and Computation, 8(1):53–60, 1979. CODEN CSSCDB. ISSN 0361-0918.

Miller:1978:PAR

[Mil78] Webb Miller. Performing armchair roundoff analyses of statistical algorithms. Communications in Statistics: Simulation and Computation, 7(3):243–255, 1978. CODEN CSSCDB. ISSN 0361-0918.

Monti:1978:NFC

[Mon78] Katherine L. Monti. A note on finding critical values for a locally optimal multivariate test statistic. Communications in Statistics: Simulation and Computation, 7(6):549–560, 1978. CODEN CSSCDB. ISSN 0361-0918.

Moses:1978:CFU

[Mos78] Lincoln E. Moses. Charts for finding upper percentage points of Student’s t in the range .01 to .00001. Communications in Statistics: Simulation and Computation, 7(5):479–490, 1978. CODEN CSS-CDB. ISSN 0361-0918.

Miller:1979:PST

[MQ79] F. L. Miller, Jr. and C. P. Quesenberry. Power studies of tests for uniformity, II. Communications in Statistics: Simulation and Computation, 8(3):271–290, 1979. CODEN CSSCDB. ISSN 0361-0918.

Murphy:1976:CST

[Mur76] B. P. Murphy. Comparison of some two sample means tests by simulation. Communications in Statistics: Simulation and Computation, 5(1):23–32, 1976. CODEN CSSCDB. ISSN 0361-0918.

Nagarsenker:1976:DDC

[Nag76] B. N. Nagarsenker. The distribution of the determinant of correlation matrix useful in principal component analysis. Communications in Statistics: Simulation and Computation, 5(1):1–13, 1976. CODEN CSSCDB. ISSN 0361-0918.

Nagaraja:1979:SRB

[Nag79] H. N. Nagaraja. Some relations between order statistics generated by different methods. Communications in Statistics: Simulation and Computation, 8(4):369–377, 1979. CODEN CSSCDB. ISSN 0361-0918.

Nichols:1979:PMS

[NG79] William G. Nichols and Jean D. Gibbons. Parameter measures of skewness. Communications in Statistics: Simulation and Computation, 8(2):161–167, 1979. CODEN CSSCDB. ISSN 0361-0918.

Neuhardt:1978:SCO

[NMC78] John B. Neuhardt and Clark A. Mount-Campbell. Selection of costoptimal 2^{k−p} fractional factorials. Communications in Statistics:

Simulation and Computation, 7(4):369–383, 1978. CODEN CSS-CDB. ISSN 0361-0918.

Narula:1977:AMS

[NW77] Subhash C. Narula and John F. Wellington. An algorithm for the minimum sum of weighted absolute errors regression. Communications in Statistics: Simulation and Computation, 6(4):341–352, 1977. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1977:EDP

[Ode77a] Robert E. Odeh. The exact distribution of Page’s L-statistic in the two-way layout. Communications in Statistics: Simulation and Computation, 6(1):49–61, 1977. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1977:ETDa

[Ode77b] Robert E. Odeh. Extended tables of the distribution of Friedman’s s-statistic in the two-way layout. Communications in Statistics: Simulation and Computation, 6(1):29–48, 1977. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1977:ETDb

[Ode77c] Robert E. Odeh. Extended tables of the distributions of rank statistics for treatment versus control in randomized block designs. Communications in Statistics: Simulation and Computation, 6(2):101– 113, 1977. CODEN CSSCDB. ISSN 0361-0918.

Owen:1977:TCL

[OH77] D. B. Owen and Tsushung A. Hua. Tables of confidence limits on the tail area of the normal distribution. Communications in Statistics: Simulation and Computation, 6(3):285–311, 1977. CODEN CSSCDB. ISSN 0361-0918.

Owen:1977:NCD

[OJ77] D. B. Owen and Faming Ju. The normal conditioned on t-distribution when the correlation is one. Communications in Statistics: Simulation and Computation, 6(2):167–179, 1977. CODEN CSSCDB. ISSN 0361-0918.

Paige:1978:NSC

[Pai78] C. C. Paige. Numerically stable computations for general univariate linear models. Communications in Statistics: Simulation and Computation, 7(5):437–453, 1978. CODEN CSSCDB. ISSN 0361-0918.

Pearson:1979:CPP

[PJB79] E. S. Pearson, N. L. Johnson, and I. W. Burr. Comparisons of the percentage points of distributions with the same first four moments, chosen from eight different systems of frequency curves. Communications in Statistics: Simulation and Computation, 8(3):191–229, 1979. CODEN CSSCDB. ISSN 0361-0918.

Patil:1977:TPP

[PKK77] S. A. Patil, J. L. Kovner, and Rudy M. King. Table of percentage points of ratios of linear combinations of order statistics of samples from exponential distributions. Communications in Statistics: Simulation and Computation, 6(2):115–136, 1977. CODEN CSSCDB. ISSN 0361-0918.

Prescott:1977:UBL

[Pre77] P. Prescott. An upper bound for any linear function of normed residuals. Communications in Statistics: Simulation and Computation, 6 (1):83–88, 1977. CODEN CSSCDB. ISSN 0361-0918.

Rosenberg:1977:SAS

[RC77] Barr Rosenberg and Daryl Carlson. A simple approximation of the sampling distribution of least absolute residuals regression estimates. Communications in Statistics: Simulation and Computation, 6(4): 421–437, 1977. CODEN CSSCDB. ISSN 0361-0918.

Richard:1976:PMS

[Ric76] Heiberger M. Richard. Posterior moments of scalar functions of a covariance matrix. Communications in Statistics: Simulation and Computation, 5(2–3):107–140, 1976. CODEN CSSCDB. ISSN 0361-0918.

Robinson:1978:C

[Rob78] Enders A. Robinson. Comments. Communications in Statistics: Simulation and Computation, 7(1):88–89, 1978. CODEN CSSCDB. ISSN 0361-0918.

Rustagi:1978:OSO

[Rus78] Jagdish S. Rustagi. Optimization in statistics: an overview. Communications in Statistics: Simulation and Computation, 7(4):303–307, 1978. CODEN CSSCDB. ISSN 0361-0918.

Schmeiser:1977:QEG

[SD77] Bruce W. Schmeiser and Stuart Jay Deutsch. Quantile estimation from grouped data: the cell midpoint. Communications in Statistics: Simulation and Computation, 6(3):221–234, 1977. CODEN CSSCDB. ISSN 0361-0918.

Sposito:1977:UAE

[SHC77] V. A. Sposito, M. L. Hand, and G. F. Mc Cormick. Using an approximate L_{1} estimator. Communications in Statistics: Simulation and Computation, 6(3):263–268, 1977. CODEN CSSCDB. ISSN 0361-0918.

Sivazlian:1978:MLO

[Siv78] B. D. Sivazlian. Maximum likelihood and optimum location of observation sites in geodetic and spatial surveying. Communications in Statistics: Simulation and Computation, 7(4):385–398, 1978. CO-DEN CSSCDB. ISSN 0361-0918.

Smit:1976:ECT

[Smi76] C. F. Smit. An empirical comparison of two types of rank tests for the C-sample problem in the univariate and bivariate cases. Communications in Statistics: Simulation and Computation, 5(2–3):93–106, 1976. CODEN CSSCDB. ISSN 0361-0918.

Smit:1979:ECS

[Smi79] C. F. Smit. An empirical comparison of several tests for parallelism of regression lines. Communications in Statistics: Simulation and Computation, 8(1):61–74, 1979. CODEN CSSCDB. ISSN 0361-0918.

Smith:1979:EAD

[SRMS79] Paul J. Smith, Donald S. Rae, Ronald W. Manderscheid, and Sam Silbergeld. Exact and approximate distributions of the chi-square statistic for equiprobability. Communications in Statistics: Simulation and Computation, 8(2):131–149, 1979. CODEN CSSCDB. ISSN 0361-0918.

Subrahmaniam:1979:TSO

[Sub79] Kathleen Subrahmaniam. Tests of significance for the odds ratio in a 2 × 2 table based on the ratio of two independent F -variates. Communications in Statistics: Simulation and Computation, 8(3): 245–255, 1979. CODEN CSSCDB. ISSN 0361-0918.

Sundberg:1976:IMS

[Sun76] Rolf Sundberg. An iterative method for solution of the likelihood equations for incomplete data from exponential families. Communications in Statistics: Simulation and Computation, 5(1):55–64, 1976. CODEN CSSCDB. ISSN 0361-0918.

Tarter:1978:NNP

[TM78] Michael E. Tarter and John S. Marshall. A new nonparametric procedure designed for simulation studies. Communications in Statistics: Simulation and Computation, 7(3):283–293, 1978. CODEN CSS-CDB. ISSN 0361-0918.

Tukey:1978:C

[Tuk78] John W. Tukey. Comment. Communications in Statistics: Simulation and Computation, 7(1):79–84, 1978. CODEN CSSCDB. ISSN 0361-0918.

Tan:1978:ACN

[TW78] W. Y. Tan and S. P. Wong. On approximating the central and noncentral multivariate gamma distributions. Communications in Statistics: Simulation and Computation, 7(3):227–242, 1978. CO-DEN CSSCDB. ISSN 0361-0918.

Ury:1977:CSA

[Ury77] Hans K. Ury. A comparison of some approximations to the Wilcoxon– Mann–Whitney distribution. Communications in Statistics: Simulation and Computation, 6(2):181–197, 1977. CODEN CSSCDB. ISSN 0361-0918.

Vishnuvajjala:1978:OPW

[Vis78] Rama Lakshmi Vishnuvajjala. On an optimization problem for a Wiener process. Communications in Statistics: Simulation and Computation, 7(4):399–415, 1978. CODEN CSSCDB. ISSN 0361-0918.

Velleman:1976:EIS

[VW76] Paul Velleman and Roy E. Welsch. On evaluating interactive statistical program packages. Communications in Statistics: Simulation and Computation, 5(4):197–205, 1976. CODEN CSSCDB. ISSN 0361-0918.

Weinert:1978:SMO

[Wei78] Howard L. Weinert. Statistical methods in optimal curve fitting. Communications in Statistics: Simulation and Computation, 7(4): 417–435, 1978. CODEN CSSCDB. ISSN 0361-0918.

Woodward:1978:NAM

[WG78] Wayne A. Woodward and H. L. Gray. New ARMA models for Wolfer’s sunspot data. Communications in Statistics: Simulation and Computation, 7(1):97–116, 1978. CODEN CSSCDB. ISSN 0361-

Williams:1978:PMC

[Wil78] George W. Williams. A pilot Monte Carlo study of two sequential estimation procedures based on generalized U-statistics. Communications in Statistics: Simulation and Computation, 7(2):129–149, 1978. CODEN CSSCDB. ISSN 0361-0918.

Wingo:1976:MTB

[Win76] Dallas R. Wingo. Moving truncations barrier-function methods for estimation in three-parameter lognormal models. Communications in Statistics: Simulation and Computation, 5(1):65–80, 1976. CODEN CSSCDB. ISSN 0361-0918.

Winterbon:1978:DPJ

[Win78] K. B. Winterbon. Determining parameters of the Johnson s_{u}. Communications in Statistics: Simulation and Computation, 7(3):223– 226, 1978. CODEN CSSCDB. ISSN 0361-0918.

Woodward:1977:SCM

[Woo77] Wayne A. Woodward. Subroutines for computing minimum variance unbiased estimators of functions of the parameters in the normal and gamma distributions. Communications in Statistics: Simulation and Computation, 6(1):63–73, 1977. CODEN CSSCDB. ISSN 0361-0918.

Woods:1977:UFS

[WP77] Jimmie D. Woods and Harry O. Posten. The use of Fourier series in the evaluation of probability distribution functions. Communications in Statistics: Simulation and Computation, 6(3):201–219, 1977. CODEN CSSCDB. ISSN 0361-0918.